# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ II $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$$\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\mathbf{\Theta}$
Observe that the bounds on $n$ endow $(\mathbf{\Theta}_{\rm ms})_{1}$ with linearly independent rows. In Case I, $(\mathbf{\Theta}_{\rm ms})_{1}\in\mathbb{C}^{(r+1)\times n}$ and $n\geq r+1$. Case II has $(\mathbf{\Theta}_{\rm ms})_{1}\in\mathbb{C}^{r\times n}$ and $n\geq r$. Therefore, Eq. 32 says $(\mathbf{\Theta}_{\rm ms})_{1}$ has full row rank.
Consequently, $\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}$ is the unique representation of $\boldsymbol{\mathbf{1}}_{n}^{H}$ in the row space of $(\mathbf{\Theta}_{\rm ms})_{1}$. Additionally, $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ i.e., $1$ is a node of $(\mathbf{\Theta}_{\rm ms})_{1}$. So, $\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}$ must be zero at all indices except that corresponding to the node at 1. Hence,
Now, Proposition D.18 says $\mathcal{R}\left((\mathbf{C}_{\rm ms})_{\not\owns 1}\right)=\mathcal{R}\left(\mathbf{C}_{\not\owns 1}\right)$. This gives:
$\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\not\owns 1}=\boldsymbol{\mathbf{0}}.$
Since we are not dealing with Case III, Lemma D.16 gives $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. So, Proposition D.18 also says that $\mathcal{R}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{C}_{\not\owns 1})$. Thus, we get,
$\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}~{}=~{}\boldsymbol{\mathbf{0}},$
which contradicts Eq. 51.
Putting together the above results gives a deceptively short proof for the non-equivalence of mean-subtracted DMD and DFT in the over-sampled regime.
Proof D.24 (Proof of Theorem 4.5).
We can preface the justification by observing that the conditions of Proposition C.3 are met, thereby endowing the matrix $\tilde{\mathbf{C}}$ with linearly independent columns.
The proof consists of two branches:
  1. 1.
    $1\in\sigma(\mathbf{\Theta}_{\rm ms}):$ Since all the conditions for Theorem D.22 are met, there is no DMD-DFT equivalence.
  2. 2.
    $1\notin\sigma(\mathbf{\Theta}_{\rm ms}):$ Since $n\geq r+1$, by Theorem D.20, there is no DMD-DFT equivalence.
D.5 Exploring the under and just-sampled regimes
Proposition D.25.
Suppose (3.8) and (3.10) are true. If the columns of $\tilde{\mathbf{C}}$ are linearly independent and $1$ is a node of $\mathbf{\Theta}_{\rm ms}$, then, $\mathcal{N}(\mathbf{C}_{\rm ms})$ is the one-dimensional subspace spanned by $\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}$.
[…]
Proof D.26.
@@ D.5 Exploring the under and just-sampled regimes
$\underbrace{\mathbf{Z}_{\rm ms}}\frac{\mathbf{1}_{n+1}}{n+1}=\boldsymbol{\mathbf{0}}~{}\iff~{}\mathbf{C}_{\rm ms}\underbrace{\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}}=\underbrace{\boldsymbol{\mathbf{0}}}~{}\implies~{}\mathcal{N}(\mathbf{C}_{\rm ms})\supseteq\mathcal{R}(\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}).$
We will show that the set containment proceeds both ways, and the condition $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ helps simplify this process.
Firstly, $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ means $\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1}~{}\neq~{}\boldsymbol{\mathbf{0}}$. Therefore, $\mathcal{R}(\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1})$ is a non-trivial subspace of dimension 1. Secondly, by the contrapositive of Proposition D.14, $1$ being a node of $\mathbf{\Theta}_{\rm ms}$ ensures that $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. Consequently, we can apply Proposition D.18 to get $\mathcal{R}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{C}_{\not\owns 1})$. By the rank-nullity theorem, we then obtain:
@@
Table 4 gives us the inputs needed to compute $\dim(\mathcal{N}(\mathbf{C}_{\rm ms}))$, and the analysis splits into two branches:
  1. 1.
    $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$ (Case I): The Vandermonde matrix $\mathbf{\Theta}_{\rm ms}$ has $r+1$ rows. So, by Eq. 49, $\mathbf{C}_{\rm ms}$ has $r+1$ columns. Since $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$, $\mathbf{C}_{\not\owns 1}$ coincides with $\mathbf{C}$ which, by (3.10), inherits from $\tilde{\mathbf{C}}$ a full column rank of $r$. Hence,
@@
  2. 2.
    $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ (Case II): The matrix $\mathbf{C}_{\rm ms}$ has $r$ columns as $\mathbf{\Theta}_{\rm ms}$ has $r$ rows. Since $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ and $\tilde{\mathbf{C}}$ has full column rank, (3.10) says that $\mathbf{C}_{\not\owns 1}$ has $r-1$ linearly independent columns. Therefore, we have:
@@
Since we already know that this null space contains the span of $\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1}$, we get:
$\mathcal{N}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}).$
@@
Proof D.27 (Proof of Corollary 4.8).
There are two distinct segments to this proof, corresponding to the under-sampled ($n<r$) and the just-sampled ($n=r$) regimes. We preface the discussion by highlighting a consequence of Proposition A.4 - the matrices $\tilde{\mathbf{C}}$ and $\mathbf{C}$ both have full column rank.
  • $n<r$: The analysis of this regime splits according to “Branch 1” in Table 4. The primary focus will be to show that $\mathbf{X}_{\rm ms}$ has full column rank. DMD-DFT equivalence follows as a consequence of Lemma 3.6.
    1. 1.
      $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$ (Cases III, IV): The matrix $\mathbf{C}_{\rm ms}$ is either $\mathbf{C}$ or $\mathbf{C}_{\not\owns 1}$, both of which have full column rank. So, $(\mathbf{\Theta}_{\rm ms})_{1}$ has at least $r-1$ rows. Since it has $n$ columns, for $n\leq r-1$, it will have full column rank according to Eq. 33. Therefore, $\mathbf{X}_{\rm ms}=\mathbf{C}_{\rm ms}(\mathbf{\Theta}_{\rm ms})_{1}$ will have linearly independent columns.
    2. 2.
      $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ (Cases I, II): To prove by contradiction, suppose
@@
      According to Eq. 50, we can expand $\mathbf{X}_{\rm ms}$ to get:
@@
      By Proposition D.25, we know $\mathcal{N}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1})$. Hence, there exists a scalar $\beta$ such that:
@@
      Table 4 says $\mathbf{\Theta}_{\rm ms}$ has $n+1$ columns and at least $r$ rows. Since $n+1\leq r$, Eq. 33 says $\mathbf{\Theta}_{\rm ms}$ has full column rank. Using this in the above expression, we have $\beta=0$ and $\boldsymbol{\mathbf{z}}=\boldsymbol{\mathbf{0}}$, which contradicts our assumption that $\boldsymbol{\mathbf{z}}\neq\boldsymbol{\mathbf{0}}$. Hence, $\mathbf{X}_{\rm ms}$ has full column rank.
@@
  • $n~{}=~{}r:$ Just-sampling necessitates a case-by-case handling of the scenarios described in Table 4.
    1. 1.
      Case I: The arguments are very similar to the proof for $n<r$. Suppose there exists $\boldsymbol{\mathbf{z}}\neq\boldsymbol{\mathbf{0}}$ satisfying $\mathbf{X}_{\rm ms}\boldsymbol{\mathbf{z}}=\boldsymbol{\mathbf{0}}$. Once again, Proposition D.25, taken together with Eq. 50, says there exists $\beta$ such that
      $\mathbf{\Theta}_{\rm ms}\begin{bmatrix}\boldsymbol{\mathbf{z}}-\beta\mathbf{1}_{n}\\ -\beta\end{bmatrix}=\boldsymbol{\mathbf{0}}.$
      In Case I, $\mathbf{\Theta}_{\rm ms}$ has dimensions $r+1\times r+1$. By Eq. 32, we find that $\mathbf{\Theta}_{\rm ms}$ is invertible. Thus, $\beta=0$ and $\boldsymbol{\mathbf{z}}=\boldsymbol{\mathbf{0}}$ which is a contradiction. Hence, $\mathbf{X}_{\rm ms}$ is full column rank which ensures DMD-DFT equivalence.
    2. 2.
      Case II: We have $n=r$ and $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ along with a full column rank $\tilde{\mathbf{C}}$. Hence, by Theorem D.22, there is no DMD-DFT equivalence.
    3. 3.
      Case III: According to Lemma D.16, $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{0}}$. So, by Eq. 28, we have $\mathbf{X}_{\rm ms}=\mathbf{X}=\mathbf{C}\mathbf{\Theta}_{1}$. Equation 33 tells us that the matrix $\mathbf{\Theta}_{1}\in\mathbb{C}^{r\times r}$ has full column rank. Hence, $\mathbf{X}_{\rm ms}$ has full column rank and, therefore, we have DMD-DFT equivalence by Lemma 3.6.
    4. 4.
@@
      Case IV: The similarity of Eq. 49 to Eq. 24 lets us view the columns of $\mathbf{Z}_{\rm ms}$ as snapshots of the same dynamical system along the same trajectory starting at $\boldsymbol{\mathbf{\upsilon}}_{1}$, but generated using a potentially different set of observables. Here, in Case IV, the new dictionary lies in the non-redundant span of the $r-1$ KEFs that correspond to the eigen-values $\sigma(\mathbf{\Theta}_{\rm ms})=\{\,\lambda_{i}\,\}_{i=1}^{r}\cap\{1\}^{c}$. The initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ continues to be spectrally informative for this modified collection of observables. Furthermore, note that
@@
      • The matrix $\mathbf{C}_{\rm ms}=\mathbf{C}_{\not\owns 1}$ has full column rank. So, by Proposition A.4, the new dictionary is also Koopman invariant- a property that, by definition, guarantees the pertinent inputs to DMD are linearly consistent.
      • The mean-subtracted time series $\mathbf{Z}_{\rm ms}$ has zero temporal mean.
@@
      Applying Theorem 4.5, we see that there is no DMD-DFT equivalence.
D.6 Delay embedding can generate a Koopman invariant dictionary

### v6

@@ E.5 Exploring the under and just-sampled regimes
Proof
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{Z}_{\rm ms}}\frac{\mathbf{1}_{n+1}}{n+1}=\boldsymbol{\mathbf{0}}~{}\iff~{}\mathbf{C}_{\rm ms}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{0}}}~{}\implies~{}\mathcal{N}(\mathbf{C}_{\rm ms})\supseteq\mathcal{R}(\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}).$
Now, we need to show that $\mathcal{N}(\mathbf{C}_{\rm ms})$ is in-turn subsumed by $\mathcal{R}(\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1})$.
This is equivalent to showing that $\mathcal{N}(\mathbf{C}_{\rm ms})$ is a one dimensional subspace. Such an equivalence arises from $1$ being a node of $\mathbf{\Theta}_{\rm ms}$, a property that ensures $\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1}~{}\neq~{}\boldsymbol{\mathbf{0}}$. Hence, $\mathcal{R}(\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1})$ is a non-trivial subspace of dimension 1. Since $\mathcal{N}(\mathbf{C}_{\rm ms})\supseteq\mathcal{R}(\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1})$, we need only show that $\mathcal{N}(\mathbf{C}_{\rm ms})$ is one-dimensional to turn the set containment into an equality.
Hence, we quantify the dimension of $\mathcal{N}(\mathbf{C}_{\rm ms})$, through the more tractable $\mathcal{R}(\mathbf{C}_{\not\owns 1})$. Once again, $1$ being a node of $\mathbf{\Theta}_{\rm ms}$ means we are either in Case I or II of Table 4. This lets us apply Lemma E.11 to infer that $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. Consequently, we can apply Proposition E.12 to get $\mathcal{R}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{C}_{\not\owns 1})$. By the rank-nullity theorem, we then obtain:
@@
Since $\dim(\mathcal{R}(\mathbf{C}_{\not\owns 1}))$ is always one less than the number of columns in $\mathbf{C}_{\rm ms}$, we have the desired conclusion. We can compute $\dim(\mathcal{R}(\mathbf{C}_{\not\owns 1}))$ using Table 4. Since $1\in\sigma(\mathbf{\Theta}_{\rm ms})$, we need only consider Cases I and II. We preface these considerations by noting that the spectrally informative initial condition ensures $\mathbf{C}$ inherits, from $\tilde{\mathbf{C}}$, a full column rank of $r$.
Case I
Here, $1$ is not a Koopman eigenvalue i.e., $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$ and the Vandermonde matrix $\mathbf{\Theta}_{\rm ms}$ has $r+1$ rows.
According to Eq. 63, $\mathbf{\Theta}_{\rm ms}$ can be pre-multiplied by $\mathbf{C}_{\rm ms}$. So, $\mathbf{C}_{\rm ms}$ must have $r+1$ columns. Since $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$, Definition E.6 says $\mathbf{C}_{\not\owns 1}$ coincides with $\mathbf{C}$ which has a full column rank of $r$. Hence,
@@
Case II
In this scenario, $1$ is a Koopman eigenvalue i.e., $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ and $\mathbf{\Theta}_{\rm ms}$ has $r$ rows.
As before, we can draw upon Eq. 63 to deduce that $\mathbf{C}_{\rm ms}$ has $r$ columns. Since $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$, Definition E.6 says that $\mathbf{C}_{\not\owns 1}$ is obtained by excluding one of the $r$ columns in $\mathbf{C}$. Hence, the full column rank of $\mathbf{C}$ translates to $\mathbf{C}_{\not\owns 1}$ possessing $r-1$ linearly independent columns. Therefore, we have:
@@
[…]
@@
E.5.1 Proof of Corollary 4.8
There are two distinct segments to this proof, corresponding to the under-sampled ($n<r$) and the just-sampled ($n=r$) regimes. We preface the discussion by highlighting a consequence of Proposition B.3 and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ being spectrally informative - the matrices $\tilde{\mathbf{C}}$ and $\mathbf{C}$ both have full column rank.
Under-sampling $(n<r)$
In this regime, we will establish the equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT using Lemma 3.9. So, we will focus on showing that $\mathbf{X}_{\rm ms}$ has full column rank.
The pertinent analysis splits according to the efficacy of mean-subtraction, as outlined in the second column of Table 4.
Cases I and II
According to Table 4, $1$ is a node of the Vandermonde matrix $\mathbf{\Theta}_{\rm ms}$ i.e., $1\in\sigma(\mathbf{\Theta}_{\rm ms})$.
Proof by contradiction. Suppose $\mathbf{X}_{\rm ms}$ does not have full column rank i.e.,
@@
Then, the matrix $\mathbf{\Theta}_{\rm ms}$ has a non-trivial nullspace. To see this, we first use Eq. 64 to expand $\mathbf{X}_{\rm ms}$ thus:
@@
By Proposition E.15, we know $\mathcal{N}(\mathbf{C}_{\rm ms})=\mathcal{R}(\mathbf{\Theta}_{\rm ms}\boldsymbol{\mathbf{1}}_{n+1})$. Hence, there exists a non-zero scalar $\beta$ such that:
@@
However, under-sampling ensures that $\mathbf{\Theta}_{\rm ms}$ has full column rank, which is a contradiction. This can be deduced from Table 4 which says $\mathbf{\Theta}_{\rm ms}$ has $n+1$ columns and $r+1$ ($r$) rows in Case I (Case II). Since under-sampling means $n+1\leq r$, there are at least as many rows in $\mathbf{\Theta}_{\rm ms}$ as columns. So, we can apply Eq. 45 to deduce that $\mathbf{\Theta}_{\rm ms}$ has full column rank.
Thus, $\mathbf{X}_{\rm ms}$ must have full column rank.
Cases III and IV
Now, the number $1$ is no longer a node of $\mathbf{\Theta}_{\rm ms}$ i.e., $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$.
The matrix $\mathbf{C}_{\rm ms}$ has full column rank. According to Table 4, the matrix $\mathbf{C}_{\rm ms}$ is $\mathbf{C}$ (Case III) or $\mathbf{C}_{\not\owns 1}$ (Case IV). Since both matrices have full column rank, $\mathbf{C}_{\rm ms}$ will also have linearly independent columns.
Similarly, the matrix $(\mathbf{\Theta}_{\rm ms})_{1}$ also has full column rank. To begin with, we have just seen that $\mathbf{C}_{\rm ms}$ possesses at least $r-1$ columns. So, by Eq. 64, $(\mathbf{\Theta}_{\rm ms})_{1}$ has at least $r-1$ rows. Since it has $n$ columns and under-sampling means $n\leq r-1$, Eq. 45 says that $(\mathbf{\Theta}_{\rm ms})_{1}$ has full column rank.
Therefore, $\mathbf{X}_{\rm ms}$, being the product of these two matrices, inherits linearly independent columns.
Just-sampling $(n~{}=~{}r)$
@@
Just-sampling necessitates a case-by-case handling of the scenarios described in Table 4.
Case I
Here, $1$ is a node of the Vandermonde matrix $\mathbf{\Theta}_{\rm ms}$ but not a Koopman eigenvalue.
The latter means that we need to show $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ is equivalent to a DFT.
To this end, we can adopt the same line of reasoning used in the under-sampled regime, with appropriate modifications in showing that $\mathbf{\Theta}_{\rm ms}$ has full column rank. By Eqs. 11 and 63, $\mathbf{\Theta}_{\rm ms}$ has $n+1$ columns. Additionally, this being Case I, Table 4 tells us that $\mathbf{\Theta}_{\rm ms}$ has $r+1$ distinct nodes, which translates to $r+1$ rows. Since just-sampling means $n~{}=~{}r$, $\mathbf{\Theta}_{\rm ms}$ still has at-least as many rows as columns. So, we can continue to rely on Eq. 45 to deduce that $\mathbf{\Theta}_{\rm ms}$ has linearly independent columns.
Case II
According to Table 4, $1$ is a node of $\mathbf{\Theta}_{\rm ms}$ and is also a Koopman eigenvalue.
So, we need to establish non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT.
This can be achieved by drawing on Theorem 4.5, even though we are not in the over-sampled regime. Firstly, note that all the conditions required by Theorem 4.5, except for over-sampling, are met. We have a dictionary in the non-redundant span of $r$ distinct KEFs and a spectrally informative initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$. Furthermore, the Koopman invariance of $\boldsymbol{\mathbf{\psi}}$ ensures that $(\mathbf{X},\mathbf{Y})$ is linearly consistent. Now, Remark E.14 tells us that here in Case II, Theorem 4.5 holds under the weaker condition of well-sampling ($n\geq r$). Since just-sampling ($n=r$) implies well-sampling, we can apply Theorem 4.5 to infer non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT.
Case III
According to Table 4, $1$ is neither a node of $\mathbf{\Theta}_{\rm ms}$ nor a Koopman eigenvalue.
So, we need to establish the equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT.
Once again, we can re-use the reasoning from Case III in the under-sampled regime, with relevant modifications to show that $(\mathbf{\Theta}_{\rm ms})_{1}$ has full column rank. Firstly, by Table 4, we have $\mathbf{C}_{\rm ms}~{}=~{}\mathbf{C}$. This means that $\mathbf{C}_{\rm ms}$ has $r$ columns. So, by Eqs. 11, 64, and 19, $(\mathbf{\Theta}_{\rm ms})_{1}$ has $r$ rows and $n$ columns. Since just-sampling means $n~{}=~{}r$, the matrix $(\mathbf{\Theta}_{\rm ms})_{1}$ continues to possess at-least as many rows as columns. Consequently, Eq. 45 tells us that $(\mathbf{\Theta}_{\rm ms})_{1}$ has full column rank.
Case IV
Here, Table 4 tells us that $1$ is not a node of $\mathbf{\Theta}_{\rm ms}$, but is a Koopman eigenvalue.
So, we need to prove that $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ is not equivalent to a DFT.
@@
This can be achieved using Theorem 4.5, despite only being just-sampled, with an appropriate “change of coordinates”. In particular, the similarity of Eq. 63 to Eq. 36 lets us view the columns of $\mathbf{Z}_{\rm ms}$ as snapshots of the same dynamical system along the same trajectory starting at $\boldsymbol{\mathbf{\upsilon}}_{1}$, but generated using a different set of observables. The new dictionary lies in the non-redundant span of the $r-1$ KEFs that correspond to the eigen-values $\{\,\lambda_{i}\,\}_{i=1}^{r}\cap\{1\}^{c}$. The initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ continues to be spectrally informative for this modified collection of observables. Furthermore, note that
@@
  • The matrix $\mathbf{C}_{\rm ms}=\mathbf{C}_{\not\owns 1}$ has full column rank. So, by Proposition B.3, the new dictionary is also Koopman invariant- a property that, by definition, guarantees the pertinent inputs to DMD are linearly consistent.
@@
Applying Theorem 4.5, we find that evaluating Algorithm 2 using $\mathbf{Z}_{\rm ms}$ as the input is not equivalent to performing a temporal DFT. However, the mean-subtracted time series $\mathbf{Z}_{\rm ms}$ has zero temporal mean. So, the output of Algorithm 2 remains the same regardless of whether we use $\mathbf{Z}$ or $\mathbf{Z}_{\rm ms}$ as the input. Therefore, when we use $\mathbf{Z}$ as the input to Algorithm 2 i.e., when we perform $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$, it is not equivalent to a DFT.
[…]
E.6 Delay embedding can generate a Koopman invariant dictionary